In isosceles triangle ABC with side BC as the base, we are given AB = 3x − 2 and AC = 4x − 5. Which is x?
7
6
5
3

Answer by umail08(87)

In this triangle side AB = AC
3x- 2 = 4x – 5
3x – 2 + 2 = 4x – 5 +2
3x = 4x – 3
3x – 4x = 4x – 4x – 3
-x = -3
x = 3